QUE// 1(a) ZIPBIT common stock is selling at a P/E of 10 times trailing earnings . The stock price is $23.50. what were the firms earnings per share .

   1(b) PEGCOR has a P/E ratio of 1.5 . Earnings per share are $ 2.00 and the expressed EPS five years from today is $ 3.22 . Calculate the PEG ratio.
Answer #1

1(a)

Market price common stock = $ 23.50

P/E = Market price of common stock / EPS

EPS = Market price of common stock / P/E

= $ 23.50 / 10

= $ 2.35

1(b)

P/E = 1.5

EPS = 2

EPS (in 5 years) = $ 3.22

years = 5

EPS (Today) * (1+ growth rate)5 = EPS (in 5 years)

2 * (1+g)5 = 3.22

1+ g = (3.22 /2 ) 1/5

Growth rate = 9.993032%

Growth rate*100 = 9.993032

PEG ratio = 1.5 / 9.993032

= 0.150105